OpenStudy (anonymous):

What is the length of the transverse axis? (y-2)^2/16-(x+1)^2/144=1

OpenStudy (asnaseer):

in a hyperbola of the form:\[\frac{(x-x_c)^2}{a^2}-\frac{(y-y_c)^2}{b^2}=1\]I believe the length of the transverse axis is 2a.

OpenStudy (anonymous):

Thank you!

OpenStudy (asnaseer):

yw

OpenStudy (asnaseer):

you might find this helpful: http://www.purplemath.com/modules/hyperbola.htm

OpenStudy (mertsj):

a^2=16 so a = 4 2a = lengths of transverse axis. In this case, 8